### 03 Mar 2026 v1

# Required-edge Cycle Cover Problem: an ASP-Completeness Framework for Graph Problems and Puzzles

Abstract
Proving the NP-completeness of pencil-and-paper puzzles typically relies on reductions from combinatorial problems such as the satisfiability problem (SAT). Although the properties of these problems are well studied, their purely combinatorial nature often does not align well with the geometric constraints of puzzles. In this paper, we introduce the Required-edge Cycle Cover Problem (RCCP)—a variant of the Cycle Cover Problem (CCP) on mixed graphs. CCP on mixed graphs was studied by Seta (2002) to establish the ASP-completeness (i.e., NP-completeness under parsimonious reductions) of the puzzle Kakuro (a.k.a. Cross Sum), and is known to be ASP-complete under certain conditions. We prove the ASP-completeness of RCCP under certain conditions, and strengthen known ASP-completeness results of CCP on mixed graphs as a corollary. Using these results, we resolve the ASP-completeness of Constraint Graph Satisfiability (CGS) in a certain case, addressing an open problem posed by the MIT Hardness Group (2024). We also show that Kakuro remains ASP-complete even when the available digit set is $\{1,2,3\}$, consequently completing its complexity classification regarding the maximum available digit and the maximum lengths of contiguous blank cells. It strengthens previously known results of Seta (2002) and Ruepp and Holzer (2010). Furthermore, we introduce a flow model equivalent to the constrained RCCP; this model allows gadgets to be tiled densely on a rectangular grid, which enables us to reduce RCCP to various pencil-and-paper puzzles in a parsimonious manner. By applying this framework, we prove the ASP-completeness of several puzzles, including Chocona, Four Cells, Hinge, and Shimaguni, and strengthen existing NP-completeness results for Choco Banana and Five Cells to ASP-completeness.
Keywords ASP-completeness, pencil-and-paper puzzles, cycle cover, Constraint Graph Satisfiability, Kakuro, Choco Banana, Five Cells
1  Introduction
In the study of pencil-and-paper puzzles, a central question is whether a given puzzle is NP-complete. Such NP-completeness results are typically established via reductions from variants of Boolean satisfiability (SAT). Although puzzles are often defined on rectangular grids and thus subject to strict geometric constraints, the SAT variants used in reductions are usually purely combinatorial. Even when planarity is considered, these combinatorial constraints are generally less restrictive than the geometric ones arising in actual puzzles. By contrast, puzzles often exhibit richer inherent structures than SAT, such as area constraints, arithmetic constraints, or shape constraints. While this richness can make analysis more difficult, it also enables the natural expression of diverse and complex constraints beyond standard Boolean logic.
Furthermore, in pencil-and-paper puzzles, the uniqueness of solutions is a fundamental requirement. To characterize the computational hardness of verifying uniqueness, the notion of ASP-completeness is used. A problem is ASP-complete if it is NP-complete under parsimonious reductions, that is, reductions that preserve the number of solutions. ASP-completeness is naturally tied to counting problems: if a problem is ASP-complete, then by definition, its counting version is #P-complete. Furthermore, for any fixed $k$, given an instance and $k$ solutions, it is NP-complete to decide whether there exists a distinct solution [YatoSeta03ASP]. (We use “ASP-completeness” for both the decision and counting versions.) While many well-known SAT variants are ASP-complete (e.g., 3-SAT [Valiant1979Permanent, YatoSeta03ASP] and Planar 3-SAT [Hunt98PlanarCounting]), constructing puzzle gadgets that strictly preserve the number of solutions remains challenging.
General frameworks for establishing ASP-completeness of puzzles have been proposed in prior research. For instance, the framework based on the Hamiltonian Cycle Problem on grid graphs is introduced in [Tang22Loop] and is well-suited for grid-based puzzles. This framework is later refined in [MITHardness24Hamiltonicity], where the ASP-completeness of various loop puzzles is established. However, it relies on global constraints to enforce Hamiltonicity, which limits its applicability to puzzles lacking such global loop structures.
Another framework is the Constraint Graph Satisfiability problem (CGS) introduced in [HearnDemaine09Puzzles, Section 5.1.3], which is a variant of the graph orientation problem defined on constraint graphs. In essence, a constraint graph is a graph with edge weights $1$ or $2$, where the vertices are of three types: and, or, and maj (majority). While planar CGS with only and and or vertices is known to be NP-complete [HearnDemaine09Puzzles, Theorem 5.4], extending this result to ASP-completeness is nontrivial. This is because the known reduction relies on a red-blue conversion gadget ([HearnDemaine09Puzzles, Figure 2.4]) that is not parsimonious, and constructing a parsimonious version is conjectured to be impossible [MITHardness24GeneralizedSAT]. To address this distinction, recent work [MITHardness24GeneralizedSAT] formalized two variants of CGS: arbitrary edge weights, where an edge may have different weights at its two endpoints, and the standard matching edge weights, where an edge must have the same weight at both endpoints; the former is essentially equivalent to a red-blue conversion. Regarding the restriction of vertex types, the complexity classification for CGS with arbitrary edge weights is now fully settled [MITHardness24GeneralizedSAT]; in particular, the case with and, or, and maj vertices is known to be ASP-complete even when the graph is planar. In contrast, the ASP-completeness of CGS with matching edge weights using and, or, and maj vertices has remained open.
1.1  Our Contributions
In this paper, we propose an ASP-completeness framework designed to exploit the rich structural properties of puzzles, and demonstrate its versatility by applying it to several puzzle types. Specifically, our framework is based on a variant of the Cycle Cover Problem (CCP), which asks whether a given graph can be covered by vertex-disjoint cycles. CCP on mixed graphs (i.e., graphs with both directed and undirected edges) was originally introduced to analyze the ASP-completeness of Kakuro (Cross Sum) puzzles, and was shown to be ASP-complete even when restricted to planar max-degree-$3$ graphs [Seta02CrossSum]. First, we introduce required edges—edges that must be included in the cycle cover—into CCP on mixed graphs, and define this problem as the Required-edge Cycle Cover Problem (RCCP). We focus on a subclass of instances on planar degree-$3$ bipartite graphs where each vertex is incident to exactly one required edge; we refer to this problem as Restricted RCCP. We prove that Restricted RCCP is ASP-complete. Moreover, by constructing a parsimonious gadget simulating a required edge, we obtain a corollary for CCP restricted to planar degree-$3$ bipartite mixed graphs where each vertex is incident to at least one directed edge. We denote this variant as Restricted CCP and conclude that it is also ASP-complete.
Second, we establish the ASP-completeness of several problems via reductions from Restricted RCCP/CCP. Crucial to these reductions is the exclusive structure of Restricted RCCP/CCP: since every vertex is of degree $3$, fixing a cycle cover enforces a strict local constraint where each vertex must have exactly one incoming edge, one outgoing edge, and one unused edge. From the edge perspective, each edge can take one of three states: being traversed in either direction or remaining unused (with directed edges limited to two states). Since each consistent assignment of these states corresponds to a unique cycle cover, this structure allows us to simulate Restricted RCCP/CCP using simple vertex and edge gadgets. Furthermore, the bipartiteness of the graphs permits the use of asymmetric edge gadgets and facilitates the geometric embedding of these gadgets into the rectangular grids typical of puzzles.
Applying this strategy, we prove the ASP-completeness of CGS with matching edge weights using and, or, and maj vertices, thereby resolving the open problem mentioned above [MITHardness24GeneralizedSAT]. We also strengthen the known ASP-completeness of Kakuro [Ruepp10Kakuro, Seta02CrossSum, YatoSeta03ASP]. Specifically, we show that Kakuro remains ASP-complete even when the available digits are restricted to $\{1,2,3\}$. Combined with previous results, this completes the complexity dichotomy of Kakuro with respect to two fundamental parameters: the maximum available digit and the maximum length of consecutive blank cells.
Finally, to bridge the gap between abstract graphs and grid-based puzzles, we introduce a flow model equivalent to Restricted RCCP. This model consists of two types of vertices: sources and sinks. Conceptually, each source supplies a specified amount of flow, and each sink requires a specified amount; these constraints map naturally to puzzle rules involving regional or arithmetic (e.g., sum) constraints. Specifically, we encode a rectilinear embedding of a Restricted RCCP instance into the flow model by placing sinks at lattice points and inserting a source into each grid segment. By aligning these vertices on a rectangular grid, we can tile the puzzle board with gadgets in a systematic manner, thereby directly translating the combinatorial hardness of Restricted RCCP into a rigid geometric configuration. This framework is particularly effective for proving ASP-completeness: since the gadgets densely fill the puzzle board, they leave no empty space, thereby eliminating unintended degrees of freedom and ensuring parsimony.
Using this framework, we establish new ASP-completeness results for several puzzles, including Chocona, Four Cells, Hinge, and Shimaguni. Moreover, we strengthen existing NP-completeness results for other puzzles, specifically Choco Banana [Iwamoto2024ChocoBanana] and Five Cells [Iwamoto2022FiveCells], to ASP-completeness.
2  Preliminaries
Definition 1.
A mixed graph $G=(V,E,A)$ is a graph consisting of a set of vertices $V$, a set of undirected edges $E$, and a set of directed edges $A$. The degree of a vertex $v$ is defined as the number of undirected edges incident to $v$ plus the number of directed edges incident to $v$ (i.e., having $v$ as a head or tail).
Definition 2.
A max-degree-$k$ graph (resp. degree-$k$ graph) is a graph in which the degree of every vertex is at most $k$ (resp. exactly $k$).
A cycle (or mixed cycle) in a mixed graph is defined as an alternating sequence of vertices and edges $v_{1}c_{1}\dots v_{k}c_{k}v_{k+1}$ that satisfies the following conditions: [^1] Two cycles are considered the same if they differ only in their starting vertex (their orientations must also be the same).
    $v_{1}=v_{k+1}$,
    all $v_{i}$ and $c_{i}$ for $1\leq i\leq k$ are distinct, and
    each $c_{i}$ is either an undirected edge $\{v_{i},v_{i+1}\}\in E$ or a directed edge $(v_{i},v_{i+1})\in A$.
Based on these notions, we define the central problem of this paper.
Definition 3.
Let $G$ be a graph. A cycle cover of $G$ is a family of (vertex-disjoint) cycles such that every vertex of $G$ is contained in exactly one cycle. The Cycle Cover Problem (CCP) asks whether there exists a cycle cover of $G$.
It is a classical result that CCP on both undirected and directed graphs is solvable in polynomial time (see, e.g., [BondyMurty2008, Section 16.4] and [Huasrsch2025CycleFactors]). In contrast, CCP on mixed graphs is known to be ASP-complete under certain constraints [Seta02CrossSum].
Theorem 4 ([Seta02CrossSum]).
CCP on planar max-degree-3 mixed graphs is ASP-complete.
3  Required-edge Cycle Cover Problem
In this section, we introduce a variant of CCP with required edges.
Definition 5.
Given a mixed graph $G=(V,E,A)$ and a subset of undirected edges $R\subseteq E$, the Required-edge Cycle Cover Problem (RCCP) asks whether there exists a cycle cover that contains all edges in $R$. Edges in $R$ are called required edges.
In this paper, we mainly consider RCCP on planar degree-3 bipartite graphs in which every vertex is incident to exactly one required edge. In what follows, we refer to this variant as Restricted RCCP.
3.1  ASP-Completeness of Restricted RCCP
Theorem 6.
Restricted RCCP is ASP-complete.
Proof.We reduce from Planar Positive 1-in-3-SAT. In this problem, the incidence graph is planar, every clause consists of three positive literals, and we ask for a truth assignment where exactly one literal per clause is true. It is parsimoniously reducible from Planar 3-SAT [Hunt98PlanarCounting], which implies that the problem is ASP-complete. In Figures˜1 and 2, black and white vertices represent the bipartition of the graph, and required edges are drawn as bidirected edges.
Figure˜1 illustrates an example of the variable gadget. This chain structure can be extended to accommodate any number of connections to clause gadgets, corresponding to the degree of the variable in the formula. The gadget admits exactly two valid cycle covers; the former corresponds to assigning the variable true (Figure˜1(b)), while the latter corresponds to false (Figure˜1(c)). Figure˜2 illustrates the clause gadget. This gadget admits exactly three valid cycle covers, each of which corresponds to an assignment where exactly one literal is true. We arrange the variable and clause gadgets according to a planar embedding of the given Planar Positive 1-in-3-SAT instance, connecting them via the dangling edges indicated in the figures. This construction introduces no edge crossings, preserving the planarity of the original instance. The resulting reduction is parsimonious; thus the theorem holds. $\blacksquare$
(a) Structure
(b) True assignment
(c) False assignment
Figure 1: An example of a variable gadget and its two valid cycle covers. Each pair of open edges indicates connection to a clause gadget. While this gadget consists of three connections, it can be extended to any number, corresponding to the degree of the variable in the formula.
Figure 2: Top row: the clause gadget. Bottom row: its three valid cycle covers. Three pairs of open edges indicate connections to variable gadgets. The terminal of each connection consists of two vertices and a required edge between them. In each valid configuration, exactly two terminals of connections are covered, enforcing that exactly one variable is assigned true.
Furthermore, a required edge can be parsimoniously simulated by the gadget shown in Figure˜3, and this substitution preserves bipartiteness. Since every vertex is incident to at least one directed edge in our construction (after this modification), the following corollary also holds.
Corollary 7.
CCP on planar degree-3 bipartite mixed graphs is ASP-complete even when every vertex is incident to a directed edge.
This corollary strengthens Theorem˜4 proven in [Seta02CrossSum]. In what follows, we refer to this variant of CCP defined in Corollary˜7 as Restricted CCP.
Figure 3: Simulation of a required edge using undirected and directed edges.
Remark 8.
It is known that, on general mixed graphs, finding a cycle cover consisting of only even cycles (or containing at least one even cycle) is NP-complete [Huasrsch2025CycleFactors]. Since every cycle in a bipartite graph is even, both problems coincide with the standard CCP in our setting; thus Corollary˜7 strengthens these NP-completeness results to ASP-completeness.
3.2  Flow Model Equivalent to Restricted RCCP
In this section we present a flow model equivalent to Restricted RCCP that facilitates reduction via grid embeddings. We begin with the definition of a flow network that we will use.
Definition 9.
A flow network is a directed graph $G=(V,E)$ equipped with a supply/demand function $w\colon V\to\mathbb{Z}$. [^2] No capacity constraints are imposed.Vertices $v$ with $w(v)>0$ are called sources and have supply $w(v)$, while vertices with $w(v)<0$ are sinks and have demand w(v)-w(v). A flow is a function $f\colon E\to\mathbb{Z}_{\geq 0}$ satisfying the conservation property
$\sum_{e\in\delta^{+}(v)}f(e)-\sum_{e\in\delta^{-}(v)}f(e)=w(v)$
for all $v\in V$, where $\delta^{-}(v)$ and $\delta^{+}(v)$ denote the sets of incoming and outgoing edges of $v$, respectively.
In order to construct the model, we assume that each source has an in-degree of 0, an out-degree of $2$, and a supply of $2$. Under this assumption, an unconstrained source admits any flow distributions in $\{(0,2),(1,1),(2,0)\}$. We additionally employ three variants that restrict the admissible flow distributions as follows:
Biased
$\{(1,1),(2,0)\}$. Must send at least $1$ unit along the first edge. We say such a source is oriented toward that edge.
Exclusive
$\{(0,2),(2,0)\}$. Sends both units to exactly one of the two outgoing edges.
Fixed
$\{(1,1)\}$. Sends one unit along each outgoing edge.
Formally, we transform an instance of Restricted RCCP into a flow network as follows. First, given a planar degree-$3$ mixed graph $G$ with required edges, we rectilinearly embed it into a grid (Figure˜4(a)). Let $H$ denote the resulting grid graph containing the embedding (Figure˜4(b)). We construct a network $N$ from $H$ based on the following rules (Figure˜4(c)):
    Each lattice point of $H$ becomes a sink whose demand equals its degree in $H$.
    Each segment (including empty ones) of $H$ is subdivided by inserting a source of supply $2$; the two resulting edges are oriented toward the adjacent sinks.
The type of each inserted source is determined by the status of the corresponding segment: segments corresponding to directed edges in $G$ become biased sources oriented consistently with the edge direction; undirected edges become unconstrained sources; required edges become exclusive sources; and empty segments become fixed sources.
(a) Rectilinear embedding of $G$
(b) Grid graph $H$
(c) Flow network NN
Figure 4: An example of the flow model construction. Labels on the sources of NN indicate their types: $u$ for unconstrained, $b$ for biased (toward the arrow), $x$ for exclusive, and $f$ for fixed.
Now, we associate cycle covers of $G$ with flows in $N$ as follows. For each vertex $v$ and its corresponding sink, an edge incident to $v$ is incoming, outgoing, or unused in the cycle cover if and only if the corresponding flow value is $2$, 0, or $1$, respectively. Under this correspondence, we have the following lemma.
{restatable}
lemmalemflowmodel Let $G$ be a planar degree-$3$ mixed graph in which every vertex is incident to exactly one required edge, and let $N$ be the flow network constructed as described above. Then there is a bijection between the cycle covers of $G$ and the flows in $N$.
Proof.We first verify the forward direction. Fix a cycle cover of $G$, and consider its corresponding cycle family in $H$. Each lattice point in $H$ is either traversed by the cycle family or not.. In the former case, the sources corresponding to the incoming and outgoing edges of the cycle contribute $2$ and 0 units of flow, respectively; each source on the remaining incident edges contributes $1$ unit. In the latter case, all incident edges are unused; thus each source contributes $1$ unit of flow. Therefore, the total inflow to the sink equals its degree in $H$ in any case.
Conversely, we show that a feasible flow in $N$ induces a valid cycle cover of $G$. Fix a feasible flow in $N$. We construct a subgraph of $G$ and orient its edges based on the flow interpretation defined above. Flow conservation at intermediate lattice points ensures that each edge of $G$ has a well-defined orientation. Furthermore, due to the biased sources, this orientation is consistent with the direction of the directed edges in $G$.
The remaining task is to show that this subgraph forms a valid cycle cover. Consider a vertex $v$ in $G$ and its corresponding sink $s$ in $N$. Assume first that $s$ has degree $3$. Since $v$ has degree $3$ and is incident to exactly one required edge, one of the three sources adjacent to $s$ is an exclusive source, which supplies either 0 or $2$ units of flow. Since the flow from each source is at most $2$, the only way that three nonnegative integers, one of which is in $\{0,2\}$, can sum to $3$ is for the multiset of contributions to be $\{0,1,2\}$. The case where $s$ has degree $4$ is similar. Indeed, exactly three of its adjacent sources correspond to edges incident to $v$, and the remaining one corresponds to the empty grid segment. Since the latter is a fixed source, it sends $1$ unit of flow to $s$; thus, the total inflow to $s$ from the former three sources must be $3$. Thus, every vertex has exactly one incoming edge (corresponding to flow value $2$) and exactly one outgoing edge (corresponding to flow value 0) in the constructed subgraph, forming a valid cycle cover. The two constructions are inverses of each other. $\blacksquare$
We next introduce a restricted version of the flow model, exploiting the bipartiteness. We employ three variants of sinks, defined as follows:
T-Shaped
Degree at least $3$. Incident to exactly one exclusive source and two unconstrained or biased sources along orthogonal directions. The remaining direction, if any, is occupied by a fixed source.
L-Shaped
Degree at least $2$. Incident to a specified pair of adjacent non-fixed sources: two unconstrained, two exclusive, or two biased (one inward and one outward). Any remaining directions are occupied by fixed sources.
Blank
Incident only to fixed sources.
The T-shaped sinks correspond to vertices in the original graph, while the three types of L-shaped sinks (unconstrained, exclusive, biased) correspond to undirected, required, and directed edges, respectively (including their orientation). This restricted flow model is sufficient for the Restricted RCCP.
Lemma 10.
Let $G$ be a planar degree-$3$ bipartite mixed graph in which every vertex is incident to exactly one required edge. We can construct a flow network $N$ for $G$ using only T-Shaped, L-shaped, and blank sinks.
Proof.We first show that we can rectilinearly embed $G$ so that each edge bends at every lattice point it passes through. In particular, we ensure that at each vertex, the required edge is orthogonal to the two remaining incident edges.
First, observe the necessary parity condition. Let $V_{1}\cup V_{2}$ be the bipartition of the vertices of $G$. Assuming that we obtain a desired embedding, consider traversing an edge from a vertex in $V_{1}$ to a vertex in $V_{2}$. Suppose that the edge starts with a horizontal segment. Since horizontal and vertical segments appear alternatingly, the $i$-th segment must be horizontal for odd $i$ and vertical for even $i$; if it starts with a vertical segment, the converse holds. Considering the coordinates of the lattice points, an edge starting horizontally from an odd (resp. even) lattice point behaves consistently with an edge starting vertically from an even (resp. odd) lattice point, in the sense that they traverse lattice points of the same parity in the same direction.
Now, we construct such an embedding. Since each vertex has degree $3$, it forms a T-shape in the rectilinear embedding. Suppose that we place a vertex at an arbitrary lattice point. We align the incident segment corresponding to the required edge along one direction, and the segments corresponding to the other incident edges along the orthogonal directions. By choosing the orientation of each T-shape according to the coordinate parity, we can ensure that all required edges share one parity and all non-required edges share the other. Given this parity assignment, the actual embedding can be readily constructed by replacing each lattice point with a unit of $2\times 2$ lattice points as in Figure˜5.
(a) Vertices
(b) Straight segment
(c) Bend segments
Figure 5: $2\times 2$ units for the embedding in Lemma˜10. Edge directions are omitted for simplicity.
Finally, using this embedding, we construct the flow network $N$. Each vertex becomes a T-shaped sink, each intermediate point on an edge becomes an L-shaped sink of the corresponding type, and each empty lattice point becomes a blank sink. $\blacksquare$
In our construction of flow networks, the total supply (twice the number of edges of $H$) equals the total demand (the sum of the degrees of vertices in $H$). This leads directly to the following lemma.
Lemma 11.
Figure˜4 holds even if the conservation property is relaxed to inequalities (i.e., $\geq w(v)$ for all $v$ or $\leq w(v)$ for all $v$).
4  ASP-Completeness of Constraint Graph Satisfiability
In this section, we formally define the CGS problem and prove its ASP-completeness.
Definition 12 ([HearnDemaine09Puzzles]).
A constraint graph is a graph with vertex weights and edge weights. Each edge can be oriented in either direction. A vertex is satisfied if and only if the sum of the weights of its incoming edges is at least the vertex weight. Given a constraint graph, the Constraint Graph Satisfiability problem (CGS) asks whether there exists an orientation of the edges that satisfies all vertices.
CGS for planar graphs is called Planar CGS. Typically, all vertex weights are restricted to $2$, and all edge weights are restricted to either $1$ or $2$. Each edge is colored according to its weight: weight-$1$ edges are red, and weight-$2$ edges are blue.
We employ three standard types of vertices commonly used in constraint logic frameworks [HearnDemaine09Puzzles]:
and
Incident to one blue edge and two red edges. The blue edge can be oriented outward only if both red edges are oriented inward.
or
Incident to three blue edges. At least one of the three edges must be oriented inward.
maj
Incident to three red edges. At least two of the three edges must be oriented inward.
and: Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
or: Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
maj: Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Three types of vertices in constraint graphs and all of their valid orientations.
Furthermore, edges are often allowed to have different weights at their endpoints; that is, an edge can have weight $2$ at one endpoint and weight $1$ at the other. Problems containing such edges are referred to as having arbitrary edge weights, while the standard case is referred to as having matching edge weights [MITHardness24GeneralizedSAT]. In this work, we focus on the latter.
Theorem 13.
Planar matching edge weights CGS with and, or, and maj vertices is ASP-complete.
Proof.We construct a reduction from Restricted CCP. We first describe the construction without considering planarity. The core idea of the reduction is as follows: Each vertex of the CCP instance is colored either blue or red according to its bipartition; this coloring corresponds to the type of the associated vertex gadget. A vertex gadget has three terminals, corresponding to its three incident edges. Each terminal admits three possible states, denoted by $I$, $O$, and $U$, representing an incoming edge, an outgoing edge, and an unused edge in a cycle cover, respectively. Each vertex gadget admits exactly those assignments of terminal states in which each of $I$, $O$, and $U$ appears exactly once. An edge gadget connects a terminal of a blue vertex gadget to a terminal of a red vertex gadget. It represents a single wire that transmits the state between the two endpoints such that $I$ and $O$ are swapped, while $U$ remains unchanged. An undirected edge gadget admits all three possible states, whereas a directed edge gadget admits only two; it forbids the state corresponding to traversal against the original edge’s direction (i.e., $I$ at the tail terminal and $O$ at the head terminal). Under this construction, any valid assignment of states to the gadgets corresponds one-to-one to a cycle cover of the original instance.
We now describe the gadgets in detail. The vertex gadgets are illustrated in Figure˜7. Each terminal of a vertex gadget consists of three parallel edges labeled $I$, $O$, and $U$, representing the three possible states. As will be demonstrated, in a terminal of a blue (resp., red) vertex gadget, exactly one of these three edges must be oriented inward (resp., outward); the label of this edge determines the state of the terminal. With this convention, it is straightforward to verify that both types of vertex gadgets admit exactly those assignments of states described above.
The edge gadgets are illustrated in Figure˜8. Each gadget consists of three parallel paths connecting a blue vertex gadget to a red vertex gadget. Each end of these paths is labeled $I$, $O$, and $U$ according to the corresponding terminal of the vertex gadget. These paths connect the corresponding labels as follows: $I$ to $O$, $O$ to $I$, and $U$ to $U$. As we show below, exactly one of these three paths must be oriented from red to blue; the pair of labels of this path determines the state of the gadget. It can be verified that the undirected edge gadget (Figure˜8(a)) admits all three possible states, whereas the directed edge gadget (Figures˜8(b) and 8(c)) forbids the specified state.
(a) Blue
(b) Red
Figure 7: Two types of vertex gadgets for CGS and an example of a valid orientation of their edges. The gadgets have three terminals, and each terminal consists of three edges labeled $I,O,U$. In each terminal, the single edge oriented differently from the other two determines the state of the terminal: $(U,O,I)$ in (a) and (b), as viewed clockwise from the top.
(a) Undirected
(b) Directed (leftward)
(c) Directed (rightward)
Figure 8: Three types of edge gadgets for CGS and an example of a valid orientation of their internal edges. The hollow vertices indicate weight-1 vertices (Figure˜9). Exactly one of the three parallel paths oriented leftward determines the state of the gadget: $(I,O)$ in (a), $(O,I)$ in (b), and $(U,U)$ in (c), respectively.
Figure 9: Simulation of a weight-1 vertex using and, or, and maj vertices, showing unique valid orientations.
First, we show the validity of the blue vertex gadget. By construction, each edge gadget can orient at most one blue edge in the constraint graph outward. Summing over the three incident edge gadgets, at most three blue edges can be oriented toward a blue vertex gadget. Conversely, as a blue vertex gadget consists of three or vertices, it requires at least three incoming blue edges. Therefore, in each terminal, exactly one blue edge is oriented inward, satisfying the desired property.
Next, we show the validity of the edge gadgets and the red vertex gadget. In an edge gadget, exactly one blue edge is oriented outward; thus, at least one of the three paths must be oriented from red to blue. From the perspective of the incident red vertex gadget, in each terminal, at most two red edges can be oriented inward. Summing over the three incident edge gadgets, at most six red edges can be oriented inward. Conversely, as a red vertex gadget consists of three maj vertices, it requires at least six incoming red edges. Therefore, similarly, in each terminal exactly two red edges are oriented from an edge gadget toward the red vertex gadget (and exactly one is oriented toward the edge gadget); this satisfies the desired property.
Finally, we address planarity. The constructions of the crossing gadgets are illustrated in Figure˜10. In a red-red crossing gadget (Figure˜10(a)), no valid orientation exists if two opposite exterior edges are oriented outward. Conversely, if two adjacent exterior edges are oriented outward, the orientations of the interior edges are uniquely determined, forcing the remaining two exterior edges to be oriented inward. Hence, the orientations of the crossing red edges are consistent. The red-blue crossing gadget (Figure˜10(b)) is constructed from two red-red crossing gadgets, while the blue-blue crossing gadget (Figure˜10(c)) consists of two red-blue crossing gadgets. It is straightforward to verify that these crossing gadgets function as intended. $\blacksquare$
(a) Red-red
(b) Red-blue
(c) Blue-blue
Figure 10: Three types of crossing gadgets for CGS used to ensure planarity.
5  Applications to Puzzles
In this section, we apply the results in Section˜3 to various pencil-and-paper puzzles.
5.1  Kakuro
Kakuro [KakuroRule], also known as Cross Sum, is a well-known puzzle. The rules of Kakuro are as follows:
  1. 1.
    Given a grid, assign digits from $1$ to $9$ to each blank cell.
  2. 2.
    For each horizontal or vertical run of contiguous blank cells, the sum of the digits in the run must equal the given clue.
  3. 3.
    No digit may be repeated within a single run.
We implicitly assume that every run contains at least two blank cells and that all blank cells are orthogonally connected.
We consider a generalized version of Kakuro introduced by Seta [Seta02CrossSum]. $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is a variant of Kakuro with the following additional constraints:
    The digits allowed in each cell are restricted to $\{1,\dots,N\}$.
    Each run of contiguous blank cells has a length of between $L_{\text{min}}$ and $L_{\text{max}}$.
By definition, we assume that $N\geq L_{\text{max}}\geq L_{\text{min}}$. Under this definition, standard Kakuro is equivalent to $(9,2,9)$-Kakuro.
Observe that for any $L^{\prime}_{\text{min}}\leq L_{\text{min}}\leq L_{\text{max}}\leq L^{\prime}_{\text{max}}$, every instance of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is also a valid instance of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro. Hence, $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is trivially reducible to $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro; in particular, it is parsimonious. It follows that, for fixed $N$, the hardness of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro (or its counting version) implies the hardness of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro for all $L^{\prime}_{\text{min}}\leq L_{\text{min}}$ and $L^{\prime}_{\text{max}}\geq L_{\text{max}}$. Similarly, the tractability of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro (or its counting version) implies the tractability of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro.
In contrast, for fixed $L_{\text{min}}$ and $L_{\text{max}}$, the complexity of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro may vary with $N$. This is because increasing $N$ enlarges the set of available digits, thereby increasing the number of potential assignments to blank cells. Thus, although $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is trivially reducible to $(N^{\prime},L_{\text{min}},L_{\text{max}})$-Kakuro for $N\leq N^{\prime}$, the reduction is not parsimonious in general.
The known complexity results for Kakuro are summarized below:
    $(N,2,2)$-Kakuro is in P for all $N$ [Seta02CrossSum].
    $(N,2,5)$-Kakuro is ASP-complete for all $N\geq 7$ [Seta02CrossSum, YatoSeta03ASP].
    $(N,1,3)$-Kakuro is ASP-complete for all $N\geq 7$ [Seta02CrossSum].
    $(9,2,4)$-Kakuro is ASP-complete [Ruepp10Kakuro].
Our main result in this section is the following.
Theorem 14.
$(N,2,3)$-Kakuro is ASP-complete for all $N\geq 3$.
This result strengthens the previously known hardness results. Moreover, it is straightforward that #$(N,2,2)$-Kakuro is in FP for all $N$, since the input puzzle board must be a $2\times 2$ block. Consequently, this theorem allows us to completely classify the counting complexity of $(N,2,L_{\text{max}})$-Kakuro for all $N\geq 2$ and $L_{\text{max}}\geq 2$.
Corollary 15.
$(N,2,L_{\text{max}})$-Kakuro is in FP for $L_{\text{max}}=2$, and ASP-complete for $N\geq L_{\text{max}}\geq 3$.
Proof of Theorem˜14.We construct a reduction from Restricted CCP. Let $G=(V_{1}\cup V_{2},E,A)$ be a planar degree-3 bipartite mixed graph in which every vertex is incident to a directed edge. The key idea is to encode the state of each edge in a cycle cover of $G$ using digits assigned to the terminals of Kakuro gadgets. Specifically, an edge directed from a vertex in $V_{1}$ to a vertex in $V_{2}$ corresponds to digit $2$, an edge in the opposite direction corresponds to digit $3$, and an unused edge corresponds to digit $1$.
We construct vertex and edge gadgets. Each gadget has two types of terminals: red and yellow. A red terminal belongs to a horizontal run within the gadget and is exposed either upwards or downwards, whereas a yellow terminal belongs to a vertical run within the gadget and is exposed either leftwards or rightwards. Gadgets are connected so that a red terminal of one gadget is identified with a yellow terminal of another, without interference of blank cells. Each vertex gadget has three terminals of the same color, determined by the bipartition of the vertex: red for $V_{1}$ and yellow for $V_{2}$. Each edge gadget connects terminals of opposite colors.
Since every vertex in $G$ is incident to at least one directed edge, it suffices to construct two types of vertex gadgets per color: one incident to an incoming edge and one incident to an outgoing edge. Figure˜12 illustrates the former. The lower-left terminal represents an incoming edge and admits digits in $\{1,3\}$ in the red vertex gadget (resp., $\{1,2\}$ in the yellow vertex gadget). The latter is obtained by replacing all clues $3$ with $4$ (resp., $4$ with $3$). Note that a $90^{\circ}$ rotation exchanges terminal colors, and thereby reverses the direction of the edge.
(a) Red
(b) Yellow
Figure 12: Vertex gadgets for Kakuro (incident to a lower-left incoming edge) and their local solutions. In three terminals of each gadget, the assigned digits are different to one another, corresponding to the three incident edges of the vertex.
Vertex gadgets are connected using three types of edge gadgets: directed (red-to-yellow and yellow-to-red), and undirected. Figure˜13 illustrates the components of a directed edge gadget from red to yellow. Each component admits exactly two local solutions, forcing the terminal values to be either both $1$ or both $2$. We can extend the gadget, bend it, or invert its terminal parity. A directed edge gadget from yellow to red is obtained by swapping clue values $3$ and $4$. The components of an undirected edge gadget are illustrated in Figure˜14. Each component admits exactly three local solutions, corresponding to terminal values $1$, 22, or $3$. As with directed edges, these gadgets can be extended and bent. By appropriately configuring the vertex gadgets, we can match the parity of any terminal that admits an undirected edge; thus no parity inversion component is needed. These edge gadgets can, indeed, be placed so as not to interfere with one another or with vertex gadgets.
(a) Straight
(b) Turn
(c) Parity inversion
Figure 13: Components of a directed edge gadget for Kakuro (from red to yellow) and an example of their local solutions. The two terminals of each component admit the same digit: either $1$ or $2$. The straight component can be extended by inserting additional blank cells with appropriate clues.
(a) Straight
(b) Turn
Figure 14: Components of an undirected edge gadget for Kakuro and an example of their local solutions. The two terminals of each component admit the same digit: either 11, 22, or $3$. The straight component can be extended at the portion that has a similar structure to the straight component of a directed edge gadget.
The resulting Kakuro instance contains runs of length only $2$ and $3$, and any solution uses only digits in $\{1,2,3\}$ regardless of the availability of other digits. This completes the proof. $\blacksquare$
5.2  Choco Banana
Choco Banana [ChocoBananaRule] is a puzzle first introduced in Puzzle Communication Nikoli Vol. 176. Given a rectangular grid, the goal is to shade some cells such that every orthogonally connected block of shaded cells forms a rectangle (or a square), while no orthogonally connected block of unshaded cells forms a rectangle. A number indicates the area of the block (shaded or unshaded) containing the cell. Choco Banana is known to be NP-complete [Iwamoto2024ChocoBanana].
Before constructing the reduction, we note an observation about Choco Banana: If three cells containing the same prime numbers $p$ form an L-shaped block, they must all be unshaded. This is verified by a careful case analysis of all shading assignments, given that a shaded rectangle of area $p$ is restricted to a $1\times p$ or $p\times 1$ shape, and no unshaded block with a clue $p$ can form such a shape. (This observation is a known technique among avid Choco Banana solvers.)
Now we construct a reduction via the flow model as illustrated in Figure˜15. The configuration of unshaded cells with clues $37$ and $13$ is fixed by the observation above. Each block of unshaded cells containing the clue $37$ corresponds to a sink, while each block of undetermined cells corresponds to a source. These gadgets correctly realize the model, as shown below: Each sink requires unshaded cells equal to the difference between its current size and $37$, which equals its degree. On the other hand, each source supplies a total of two unshaded cells to the adjacent sinks (Figure˜16). Figure˜17 illustrates three variants of the sources, each of which restricts the admissible flow distributions as required.
Figure 15: Global arrangement of sources and sinks for Choco Banana. Degree-22 sink at upper left, degree-33 sinks at upper right and lower left, and degree-44 sink at lower right.
(a) Flow values $(0,2)$
(b) Flow values $(1,1)$
(c) Flow values $(2,0)$
Figure 16: Source for Choco Banana and its three local solutions, corresponding to the flow distributions.
(a) Biased (rightward)
(b) Exclusive
(c) Fixed
Figure 17: Variants of the sources for Choco Banana.
Because this reduction is parsimonious, we obtain the following result.
Theorem 16.
Choco Banana is ASP-complete.
5.3  Shimaguni
Shimaguni [ShimaguniRule] (lit. island country) is a puzzle first introduced in Puzzle Communication Nikoli Vol. 117. Given a rectangular grid partitioned into regions, the goal is to shade some cells according to the following rules: Each region must contain exactly one orthogonally connected block of shaded cells (called an “island”). A number in a region indicates the area of the island in that region. Islands in different regions cannot be orthogonally adjacent. Finally, islands in orthogonally adjacent regions must have different sizes.
We construct a reduction via the flow model as illustrated in Figure˜18. Each cross-shaped region containing a clue $9$ corresponds to a sink, while each region with a clue $7$ connecting two arms of sinks corresponds to a source. These sinks and sources function as intended, as we demonstrate below: Each source supplies a total of two unshaded cells to the two connected sinks (Figure˜19). On the other hand, each sink accepts at most as many unshaded cells as its degree from its adjacent sources, due to the difference between its area and clue value. By Lemma˜11, this suffices for the reduction. Figure˜20 illustrates three variants of sources, each of which restricts the admissible flow distributions as required.
Figure 18: Global arrangement of sources and sinks for Shimaguni. Degree-22 sink at upper left, degree-33 sinks at upper right and lower left, and degree-44 sink at lower right.
(a) Flow values $(0,2)$
(b) Flow values $(1,1)$
(c) Flow values $(2,0)$
Figure 19: Source for Shimaguni and its three local solutions. In each solution, the number of unshaded cells in the sinks corresponds to the flow distributions.
(a) Biased (rightward)
(b) Exclusive
(c) Fixed
Figure 20: Variants of the sources for Shimaguni.
Because this reduction is parsimonious, we obtain the following result.
Theorem 17.
Shimaguni is ASP-complete.
5.4  Hinge
Hinge [HingeRule] is a puzzle first introduced in Puzzle Communication Nikoli Vol. 159. Given a rectangular grid partitioned into regions, the goal is to shade cells according to the following rules: Each orthogonally connected block of shaded cells must be intersected exactly once by a single straight line segment of region boundaries (called a “hinge”). Furthermore, each such block must be symmetric with respect to its hinge. A number in a region indicates the number of shaded cells contained in that region.
We construct a reduction via the flow model as illustrated in Figure˜21. Each region with arms corresponds to a sink with degree equal to the number of arms, while each $1\times 3$ or $3\times 1$ region that connects two arms corresponds to a source. These gadgets correctly realize the model, as shown below: Each source supplies a total of two shaded cells to the two connected sinks (Figure˜22). On the other hand, each sink requires at least as many shaded cells as its degree from its adjacent sources. By Lemma˜11, this suffices for the reduction. Figure˜23 illustrates three variants of sources, each of which restricts the admissible flow distributions as required.
Figure 21: Global arrangement of sources and sinks for Hinge. Degree-22 sink at upper left, degree-33 sinks at upper right and lower left, and degree-44 sink at lower right.
(a) Flow values $(0,2)$
(b) Flow values $(1,1)$
(c) Flow values $(2,0)$
Figure 22: Source for Hinge and its three local solutions, corresponding to the flow distributions.
(a) Biased (rightward)
(b) Exclusive
(c) Fixed
Figure 23: Variants of the sources for Hinge.
These sources can be arranged so as not to interfere with one another. Because this reduction is parsimonious, we obtain the following result.
Theorem 18.
Hinge is ASP-complete.
5.5  Chocona
Chocona [ChoconaRule] is a puzzle first introduced in Puzzle Communication Nikoli Vol. 122. Given a rectangular grid partitioned into regions, the goal is to shade some cells so that every orthogonally connected block of shaded cells forms a rectangle (or a square). Each number indicates the number of shaded cells contained in that region.
We construct a reduction via the flow model as illustrated in Figure˜24. Each region with four arms corresponds to a sink, while each block of undetermined cells that connects two arms of sinks corresponds to a source. These gadgets correctly realize the model, as shown below: Each source supplies a total of two shaded cells to the two connected sinks (Figure˜25). On the other hand, each sink contains a clue value equal to its degree plus one and exactly one extra undetermined cell other than the connections to sources; thus, each sink requires at least as many shaded cells as its degree from its adjacent sources. By Lemma˜11, this suffices for the reduction. Figure˜26 illustrates three variants of the sources, each of which restricts the admissible flow distributions as required.
Figure 24: Global arrangement of sources and sinks for Chocona. Degree-22 sink at upper left, degree-33 sinks at upper right and lower left, and degree-44 sink at lower right.
(a) Flow values $(0,2)$
(b) Flow values $(1,1)$
(c) Flow values $(2,0)$
Figure 25: Source for Chocona and its three local solutions, corresponding to the flow distributions.
(a) Biased (rightward)
(b) Exclusive
(c) Fixed
Figure 26: Variants of the sources for Chocona.
Because this reduction is parsimonious, we obtain the following result.
Theorem 19.
Chocona is ASP-complete.
5.6  Five Cells
Five Cells [FiveCellsRule] is a puzzle first introduced in Puzzle Communication Nikoli Vol. 133. The goal is to partition a board into pentominoes. A number in a cell indicates how many edges of that cell belong to the pentomino boundary. Five Cells is known to be NP-complete [Iwamoto2022FiveCells].
We construct a reduction via the flow model. Figure˜27 illustrates a unit gadget. The configuration of the surrounding yellow region is uniquely determined by the boundary conditions, specifically where the gadget abuts the board border or adjacent gadgets. The central cell of the gadget corresponds to a sink, from which four wires emanate in the four cardinal directions. Conversely, the interface between two wires of adjacent unit gadgets corresponds to a source. These gadgets correctly realize the model, as shown below: The central cell requires a total of $4$ cells from the four wires to form a pentomino, while each wire supplies a specific number of cells to the central cell (Figure˜28). Specifically, a wire connected to the board border supplies exactly $1$ cell, effectively adjusting the remaining demand of the central cell to $3$ (when abutting one border) or $2$ (when abutting two borders). For the interface between two wires, the sum of their supplies is constrained to $2$; collectively, these two connected wires function as a source.
Figure 27: Unit gadget for Five Cells. The undetermined portion is divided into the central cell and four wires. The green cells indicate the connections between the central cell and the wires.
(a) Flow value 0
(b) Flow value 11
(c) Flow value 22
Figure 28: A wire of the unit gadget for Five Cells and its three local solutions. The number of green cells that is connected to the central cell correspond to flow values. The end opposite to the central cell connects to the board border or such an end of another wire.
To simulate the variants of sources, we introduce two specific types of wires (Figure˜29). An incoming wire (Figure˜29(a)) can supply either $1$ or $2$ cells, whereas an exclusive wire (Figure˜29(b)) is restricted to supplying either 0 or $2$ cells. A source with exactly one incoming wire corresponds to a biased source oriented toward the incoming side. A source with two incoming wires corresponds to a fixed source, as the only pair of flow distributions summing to $2$ is $(1,1)$. A source composed of exclusive wires corresponds to an exclusive source.
(a) Incoming
(b) Exclusive
Figure 29: Two specific types of wires for Five Cells, each restricting the local solutions and thereby the admissible flow distributions.
Because this reduction is parsimonious, we obtain the following result.
Theorem 20.
Five Cells is ASP-complete.
5.7  Four Cells
Four Cells [FourCellsRule] is a puzzle first introduced in Puzzle Communication Nikoli Vol. 132. The goal is to partition a board into tetrominoes. A number in a cell indicates how many edges of that cell belong to the tetromino boundary.
We construct a reduction based on Lemma˜10. The overall framework is analogous to the Five Cells case described in Section˜5.6. Figure˜30 illustrates the unit gadgets designed for T-shaped and L-shaped sinks. Unlike the Five Cells case, these gadgets require a checkerboard arrangement of standard and inverted orientations to ensure consistent wire connections. Regarding the flow constraints, each wire supplies a specific number of cells to the central green region (Figure˜31), whose demand is set equal to the number of incident wires. The remainder of the construction, including the wire variants (Figure˜32), follows the same principles as in the Five Cells case.
(a) T-shaped
(b) L-shaped
(c) Fixed
Figure 30: Unit gadgets for Four Cells. The undetermined portion is divided into three (T-shaped) or two (L-shaped) wires. The green cells indicate the connections between the wires.
(a) Flow value 0
(b) Flow value 11
(c) Flow value 22
Figure 31: Three local solutions of a wire. The number of cells in which the tetromino extends into the green cells corresponds to the flow value.
(a) Incoming
(b) Exclusive
Figure 32: Two specific types of wires for Four Cells, each restricting the local solutions and thereby the admissible flow distributions.
Because this reduction is parsimonious, we obtain the following result.
Theorem 21.
Four Cells is ASP-complete.
6  Conclusion and Future Work
In this paper, we established the ASP-completeness of RCCP (Theorem˜6) and CCP (Corollary˜7) for planar degree-$3$ bipartite mixed graphs. This result is tight in the sense of the maximum degree, as the counting versions of these problems on max-degree-$2$ graphs are obviously polynomial-time solvable. However, the complexity for degree-$k$ graphs with $k\geq 4$ remains open. Moreover, while the input graph in Corollary˜7 is constrained such that each vertex is incident to at least one directed edge, it remains an open question whether the problem remains ASP-complete when restricted to exactly one directed edge per vertex.
Another interesting direction is to apply the flow model to puzzles defined on other grid topologies. For instance, this model appears readily applicable to non-square grids, such as hexagonal or triangular grids. It would be worthwhile to explore what types of sinks are sufficient to simulate RCCP on such grids.

### 06 Mar 2026 v2

@@
 Another framework is the Constraint Graph Satisfiability problem (CGS) introduced in [HearnDemaine09Puzzles, Section 5.1.3], which is a variant of the graph orientation problem defined on constraint graphs.
-In essence, a constraint graph is a graph with edge weights $1$ or $2$, where the vertices are of three types: and, or, and maj (majority).
+In essence, a constraint graph is a graph with edge weights 11 or 22, where the vertices are of three types: and, or, and maj (majority).
 While planar CGS with only and and or vertices is known to be NP-complete [HearnDemaine09Puzzles, Theorem 5.4], extending this result to ASP-completeness is nontrivial.
@@
 Specifically, our framework is based on a variant of the Cycle Cover Problem (CCP), which asks whether a given graph can be covered by vertex-disjoint cycles.
-CCP on mixed graphs (i.e., graphs with both directed and undirected edges) was originally introduced to analyze the ASP-completeness of Kakuro (Cross Sum) puzzles, and was shown to be ASP-complete even when restricted to planar max-degree-$3$ graphs [Seta02CrossSum].
+CCP on mixed graphs (i.e., graphs with both directed and undirected edges) was originally introduced to analyze the ASP-completeness of Kakuro (Cross Sum) puzzles, and was shown to be ASP-complete even when restricted to planar max-degree-33 graphs [Seta02CrossSum].
 First, we introduce required edges—edges that must be included in the cycle cover—into CCP on mixed graphs, and define this problem as the Required-edge Cycle Cover Problem (RCCP).
-We focus on a subclass of instances on planar degree-$3$ bipartite graphs where each vertex is incident to exactly one required edge; we refer to this problem as Restricted RCCP.
+We focus on a subclass of instances on planar degree-33 bipartite graphs where each vertex is incident to exactly one required edge; we refer to this problem as Restricted RCCP.
 We prove that Restricted RCCP is ASP-complete.
-Moreover, by constructing a parsimonious gadget simulating a required edge, we obtain a corollary for CCP restricted to planar degree-$3$ bipartite mixed graphs where each vertex is incident to at least one directed edge.
+Moreover, by constructing a parsimonious gadget simulating a required edge, we obtain a corollary for CCP restricted to planar degree-33 bipartite mixed graphs where each vertex is incident to at least one directed edge.
 We denote this variant as Restricted CCP and conclude that it is also ASP-complete.
 Second, we establish the ASP-completeness of several problems via reductions from Restricted RCCP/CCP.
-Crucial to these reductions is the exclusive structure of Restricted RCCP/CCP: since every vertex is of degree $3$, fixing a cycle cover enforces a strict local constraint where each vertex must have exactly one incoming edge, one outgoing edge, and one unused edge.
+Crucial to these reductions is the exclusive structure of Restricted RCCP/CCP: since every vertex is of degree 33, fixing a cycle cover enforces a strict local constraint where each vertex must have exactly one incoming edge, one outgoing edge, and one unused edge.
 From the edge perspective, each edge can take one of three states: being traversed in either direction or remaining unused (with directed edges limited to two states).
@@
 Definition 1.
-A mixed graph $G=(V,E,A)$ is a graph consisting of a set of vertices $V$, a set of undirected edges $E$, and a set of directed edges $A$.
+A mixed graph $G=(V,E,A)$ is a graph consisting of a set of vertices $V$, a set of undirected edges $E$, and a set of directed edges AA.
 The degree of a vertex $v$ is defined as the number of undirected edges incident to $v$ plus the number of directed edges incident to $v$ (i.e., having $v$ as a head or tail).
@@
 for all $v\in V$, where $\delta^{-}(v)$ and $\delta^{+}(v)$ denote the sets of incoming and outgoing edges of $v$, respectively.
-In order to construct the model, we assume that each source has an in-degree of 0, an out-degree of $2$, and a supply of $2$.
+In order to construct the model, we assume that each source has an in-degree of 0, an out-degree of 22, and a supply of 22.
 Under this assumption, an unconstrained source admits any flow distributions in $\{(0,2),(1,1),(2,0)\}$.
@@
 Biased
-$\{(1,1),(2,0)\}$. Must send at least $1$ unit along the first edge. We say such a source is oriented toward that edge.
+$\{(1,1),(2,0)\}$. Must send at least 11 unit along the first edge. We say such a source is oriented toward that edge.
 Exclusive
@@
 Formally, we transform an instance of Restricted RCCP into a flow network as follows.
-First, given a planar degree-$3$ mixed graph $G$ with required edges, we rectilinearly embed it into a grid (Figure˜4(a)).
+First, given a planar degree-33 mixed graph $G$ with required edges, we rectilinearly embed it into a grid (Figure˜4(a)).
 Let $H$ denote the resulting grid graph containing the embedding (Figure˜4(b)).
-We construct a network $N$ from $H$ based on the following rules (Figure˜4(c)):
+We construct a network NN from $H$ based on the following rules (Figure˜4(c)):
     Each lattice point of $H$ becomes a sink whose demand equals its degree in $H$.
-    Each segment (including empty ones) of $H$ is subdivided by inserting a source of supply $2$; the two resulting edges are oriented toward the adjacent sinks.
+    Each segment (including empty ones) of $H$ is subdivided by inserting a source of supply 22; the two resulting edges are oriented toward the adjacent sinks.
 The type of each inserted source is determined by the status of the corresponding segment: segments corresponding to directed edges in $G$ become biased sources oriented consistently with the edge direction; undirected edges become unconstrained sources; required edges become exclusive sources; and empty segments become fixed sources.
@@
 Figure 4: An example of the flow model construction. Labels on the sources of NN indicate their types: $u$ for unconstrained, $b$ for biased (toward the arrow), $x$ for exclusive, and $f$ for fixed.
-Now, we associate cycle covers of $G$ with flows in $N$ as follows.
-For each vertex $v$ and its corresponding sink, an edge incident to $v$ is incoming, outgoing, or unused in the cycle cover if and only if the corresponding flow value is $2$, 0, or $1$, respectively.
+Now, we associate cycle covers of $G$ with flows in NN as follows.
+For each vertex $v$ and its corresponding sink, an edge incident to $v$ is incoming, outgoing, or unused in the cycle cover if and only if the corresponding flow value is 22, 0, or 11, respectively.
 Under this correspondence, we have the following lemma.
 {restatable}
-lemmalemflowmodel Let $G$ be a planar degree-$3$ mixed graph in which every vertex is incident to exactly one required edge, and let $N$ be the flow network constructed as described above.
-Then there is a bijection between the cycle covers of $G$ and the flows in $N$.
+lemmalemflowmodel Let $G$ be a planar degree-33 mixed graph in which every vertex is incident to exactly one required edge, and let NN be the flow network constructed as described above.
+Then there is a bijection between the cycle covers of $G$ and the flows in NN.
 Proof.We first verify the forward direction.
@@
 Each lattice point in $H$ is either traversed by the cycle family or not..
-In the former case, the sources corresponding to the incoming and outgoing edges of the cycle contribute $2$ and 0 units of flow, respectively; each source on the remaining incident edges contributes $1$ unit.
-In the latter case, all incident edges are unused; thus each source contributes $1$ unit of flow.
+In the former case, the sources corresponding to the incoming and outgoing edges of the cycle contribute 22 and 0 units of flow, respectively; each source on the remaining incident edges contributes 11 unit.
+In the latter case, all incident edges are unused; thus each source contributes 11 unit of flow.
 Therefore, the total inflow to the sink equals its degree in $H$ in any case.
-Conversely, we show that a feasible flow in $N$ induces a valid cycle cover of $G$.
-Fix a feasible flow in $N$.
+Conversely, we show that a feasible flow in NN induces a valid cycle cover of $G$.
+Fix a feasible flow in NN.
 We construct a subgraph of $G$ and orient its edges based on the flow interpretation defined above.
@@
 The remaining task is to show that this subgraph forms a valid cycle cover.
-Consider a vertex $v$ in $G$ and its corresponding sink $s$ in $N$.
-Assume first that $s$ has degree $3$.
-Since $v$ has degree $3$ and is incident to exactly one required edge, one of the three sources adjacent to $s$ is an exclusive source, which supplies either 0 or $2$ units of flow.
-Since the flow from each source is at most $2$, the only way that three nonnegative integers, one of which is in $\{0,2\}$, can sum to $3$ is for the multiset of contributions to be $\{0,1,2\}$.
-The case where $s$ has degree $4$ is similar.
+Consider a vertex $v$ in $G$ and its corresponding sink $s$ in NN.
+Assume first that $s$ has degree 33.
+Since $v$ has degree 33 and is incident to exactly one required edge, one of the three sources adjacent to $s$ is an exclusive source, which supplies either 0 or 22 units of flow.
+Since the flow from each source is at most 22, the only way that three nonnegative integers, one of which is in $\{0,2\}$, can sum to 33 is for the multiset of contributions to be $\{0,1,2\}$.
+The case where $s$ has degree 44 is similar.
 Indeed, exactly three of its adjacent sources correspond to edges incident to $v$, and the remaining one corresponds to the empty grid segment.
-Since the latter is a fixed source, it sends $1$ unit of flow to $s$; thus, the total inflow to $s$ from the former three sources must be $3$.
-Thus, every vertex has exactly one incoming edge (corresponding to flow value $2$) and exactly one outgoing edge (corresponding to flow value 0) in the constructed subgraph, forming a valid cycle cover.
+Since the latter is a fixed source, it sends 11 unit of flow to $s$; thus, the total inflow to $s$ from the former three sources must be 33.
+Thus, every vertex has exactly one incoming edge (corresponding to flow value 22) and exactly one outgoing edge (corresponding to flow value 0) in the constructed subgraph, forming a valid cycle cover.
 The two constructions are inverses of each other.
@@
 T-Shaped
-Degree at least $3$. Incident to exactly one exclusive source and two unconstrained or biased sources along orthogonal directions. The remaining direction, if any, is occupied by a fixed source.
+Degree at least 33. Incident to exactly one exclusive source and two unconstrained or biased sources along orthogonal directions. The remaining direction, if any, is occupied by a fixed source.
 L-Shaped
-Degree at least $2$.
+Degree at least 22.
 Incident to a specified pair of adjacent non-fixed sources: two unconstrained, two exclusive, or two biased (one inward and one outward).
@@
 Lemma 10.
-Let $G$ be a planar degree-$3$ bipartite mixed graph in which every vertex is incident to exactly one required edge.
-We can construct a flow network $N$ for $G$ using only T-Shaped, L-shaped, and blank sinks.
+Let $G$ be a planar degree-33 bipartite mixed graph in which every vertex is incident to exactly one required edge.
+We can construct a flow network NN for $G$ using only T-Shaped, L-shaped, and blank sinks.
 Proof.We first show that we can rectilinearly embed $G$ so that each edge bends at every lattice point it passes through.
@@
 Now, we construct such an embedding.
-Since each vertex has degree $3$, it forms a T-shape in the rectilinear embedding.
+Since each vertex has degree 33, it forms a T-shape in the rectilinear embedding.
 Suppose that we place a vertex at an arbitrary lattice point.
@@
 Figure 5: $2\times 2$ units for the embedding in Lemma˜10. Edge directions are omitted for simplicity.
-Finally, using this embedding, we construct the flow network $N$.
+Finally, using this embedding, we construct the flow network NN.
 Each vertex becomes a T-shaped sink, each intermediate point on an edge becomes an L-shaped sink of the corresponding type, and each empty lattice point becomes a blank sink.
@@
 CGS for planar graphs is called Planar CGS.
-Typically, all vertex weights are restricted to $2$, and all edge weights are restricted to either $1$ or $2$.
-Each edge is colored according to its weight: weight-$1$ edges are red, and weight-$2$ edges are blue.
+Typically, all vertex weights are restricted to 22, and all edge weights are restricted to either 11 or 22.
+Each edge is colored according to its weight: weight-11 edges are red, and weight-22 edges are blue.
 We employ three standard types of vertices commonly used in constraint logic frameworks [HearnDemaine09Puzzles]:
@@
 Figure 6: Three types of vertices in constraint graphs and all of their valid orientations.
-Furthermore, edges are often allowed to have different weights at their endpoints; that is, an edge can have weight $2$ at one endpoint and weight $1$ at the other.
+Furthermore, edges are often allowed to have different weights at their endpoints; that is, an edge can have weight 22 at one endpoint and weight 11 at the other.
 Problems containing such edges are referred to as having arbitrary edge weights, while the standard case is referred to as having matching edge weights [MITHardness24GeneralizedSAT].
@@
 A vertex gadget has three terminals, corresponding to its three incident edges.
-Each terminal admits three possible states, denoted by $I$, $O$, and $U$, representing an incoming edge, an outgoing edge, and an unused edge in a cycle cover, respectively.
-Each vertex gadget admits exactly those assignments of terminal states in which each of $I$, $O$, and $U$ appears exactly once.
+Each terminal admits three possible states, denoted by II, $O$, and $U$, representing an incoming edge, an outgoing edge, and an unused edge in a cycle cover, respectively.
+Each vertex gadget admits exactly those assignments of terminal states in which each of II, $O$, and $U$ appears exactly once.
 An edge gadget connects a terminal of a blue vertex gadget to a terminal of a red vertex gadget.
-It represents a single wire that transmits the state between the two endpoints such that $I$ and $O$ are swapped, while $U$ remains unchanged.
-An undirected edge gadget admits all three possible states, whereas a directed edge gadget admits only two; it forbids the state corresponding to traversal against the original edge’s direction (i.e., $I$ at the tail terminal and $O$ at the head terminal).
+It represents a single wire that transmits the state between the two endpoints such that II and $O$ are swapped, while $U$ remains unchanged.
+An undirected edge gadget admits all three possible states, whereas a directed edge gadget admits only two; it forbids the state corresponding to traversal against the original edge’s direction (i.e., II at the tail terminal and $O$ at the head terminal).
 Under this construction, any valid assignment of states to the gadgets corresponds one-to-one to a cycle cover of the original instance.
@@
 The vertex gadgets are illustrated in Figure˜7.
-Each terminal of a vertex gadget consists of three parallel edges labeled $I$, $O$, and $U$, representing the three possible states.
+Each terminal of a vertex gadget consists of three parallel edges labeled II, $O$, and $U$, representing the three possible states.
 As will be demonstrated, in a terminal of a blue (resp., red) vertex gadget, exactly one of these three edges must be oriented inward (resp., outward); the label of this edge determines the state of the terminal.
@@
 Each gadget consists of three parallel paths connecting a blue vertex gadget to a red vertex gadget.
-Each end of these paths is labeled $I$, $O$, and $U$ according to the corresponding terminal of the vertex gadget.
-These paths connect the corresponding labels as follows: $I$ to $O$, $O$ to $I$, and $U$ to $U$.
+Each end of these paths is labeled II, $O$, and $U$ according to the corresponding terminal of the vertex gadget.
+These paths connect the corresponding labels as follows: II to $O$, $O$ to II, and $U$ to $U$.
 As we show below, exactly one of these three paths must be oriented from red to blue; the pair of labels of this path determines the state of the gadget.
@@
   1. 1.
-    Given a grid, assign digits from $1$ to $9$ to each blank cell.
+    Given a grid, assign digits from 11 to 99 to each blank cell.
   2. 2.
@@
 Hence, $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is trivially reducible to $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro; in particular, it is parsimonious.
-It follows that, for fixed $N$, the hardness of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro (or its counting version) implies the hardness of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro for all $L^{\prime}_{\text{min}}\leq L_{\text{min}}$ and $L^{\prime}_{\text{max}}\geq L_{\text{max}}$.
+It follows that, for fixed NN, the hardness of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro (or its counting version) implies the hardness of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro for all $L^{\prime}_{\text{min}}\leq L_{\text{min}}$ and $L^{\prime}_{\text{max}}\geq L_{\text{max}}$.
 Similarly, the tractability of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro (or its counting version) implies the tractability of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro.
-In contrast, for fixed $L_{\text{min}}$ and $L_{\text{max}}$, the complexity of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro may vary with $N$.
-This is because increasing $N$ enlarges the set of available digits, thereby increasing the number of potential assignments to blank cells.
+In contrast, for fixed $L_{\text{min}}$ and $L_{\text{max}}$, the complexity of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro may vary with NN.
+This is because increasing NN enlarges the set of available digits, thereby increasing the number of potential assignments to blank cells.
 Thus, although $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is trivially reducible to $(N^{\prime},L_{\text{min}},L_{\text{max}})$-Kakuro for $N\leq N^{\prime}$, the reduction is not parsimonious in general.
 The known complexity results for Kakuro are summarized below:
-    $(N,2,2)$-Kakuro is in P for all $N$ [Seta02CrossSum].
+    $(N,2,2)$-Kakuro is in P for all NN [Seta02CrossSum].
     $(N,2,5)$-Kakuro is ASP-complete for all $N\geq 7$ [Seta02CrossSum, YatoSeta03ASP].
@@
 This result strengthens the previously known hardness results.
-Moreover, it is straightforward that #$(N,2,2)$-Kakuro is in FP for all $N$, since the input puzzle board must be a $2\times 2$ block.
+Moreover, it is straightforward that #$(N,2,2)$-Kakuro is in FP for all NN, since the input puzzle board must be a $2\times 2$ block.
 Consequently, this theorem allows us to completely classify the counting complexity of $(N,2,L_{\text{max}})$-Kakuro for all $N\geq 2$ and $L_{\text{max}}\geq 2$.
@@
 The key idea is to encode the state of each edge in a cycle cover of $G$ using digits assigned to the terminals of Kakuro gadgets.
-Specifically, an edge directed from a vertex in $V_{1}$ to a vertex in $V_{2}$ corresponds to digit $2$, an edge in the opposite direction corresponds to digit $3$, and an unused edge corresponds to digit $1$.
+Specifically, an edge directed from a vertex in $V_{1}$ to a vertex in $V_{2}$ corresponds to digit 22, an edge in the opposite direction corresponds to digit 33, and an unused edge corresponds to digit 11.
 We construct vertex and edge gadgets.
@@
 The lower-left terminal represents an incoming edge and admits digits in $\{1,3\}$ in the red vertex gadget (resp., $\{1,2\}$ in the yellow vertex gadget).
-The latter is obtained by replacing all clues $3$ with $4$ (resp., $4$ with $3$).
+The latter is obtained by replacing all clues 33 with 44 (resp., 44 with 33).
 Note that a $90^{\circ}$ rotation exchanges terminal colors, and thereby reverses the direction of the edge.
@@
 Figure˜13 illustrates the components of a directed edge gadget from red to yellow.
-Each component admits exactly two local solutions, forcing the terminal values to be either both $1$ or both $2$.
+Each component admits exactly two local solutions, forcing the terminal values to be either both 11 or both 22.
 We can extend the gadget, bend it, or invert its terminal parity.
-A directed edge gadget from yellow to red is obtained by swapping clue values $3$ and $4$.
+A directed edge gadget from yellow to red is obtained by swapping clue values 33 and 44.
 The components of an undirected edge gadget are illustrated in Figure˜14.
-Each component admits exactly three local solutions, corresponding to terminal values $1$, 22, or $3$.
+Each component admits exactly three local solutions, corresponding to terminal values 11, 22, or 33.
 As with directed edges, these gadgets can be extended and bent.
@@
 Figure 13: Components of a directed edge gadget for Kakuro (from red to yellow) and an example of their local solutions.
-The two terminals of each component admit the same digit: either $1$ or $2$.
+The two terminals of each component admit the same digit: either 11 or 22.
 The straight component can be extended by inserting additional blank cells with appropriate clues.
@@
 Figure 14: Components of an undirected edge gadget for Kakuro and an example of their local solutions.
-The two terminals of each component admit the same digit: either 11, 22, or $3$.
+The two terminals of each component admit the same digit: either 11, 22, or 33.
 The straight component can be extended at the portion that has a similar structure to the straight component of a directed edge gadget.
-The resulting Kakuro instance contains runs of length only $2$ and $3$, and any solution uses only digits in $\{1,2,3\}$ regardless of the availability of other digits.
+The resulting Kakuro instance contains runs of length only 22 and 33, and any solution uses only digits in $\{1,2,3\}$ regardless of the availability of other digits.
 This completes the proof.
@@
 Now we construct a reduction via the flow model as illustrated in Figure˜15.
-The configuration of unshaded cells with clues $37$ and $13$ is fixed by the observation above.
-Each block of unshaded cells containing the clue $37$ corresponds to a sink, while each block of undetermined cells corresponds to a source.
-These gadgets correctly realize the model, as shown below: Each sink requires unshaded cells equal to the difference between its current size and $37$, which equals its degree.
+The configuration of unshaded cells with clues 3737 and 1313 is fixed by the observation above.
+Each block of unshaded cells containing the clue 3737 corresponds to a sink, while each block of undetermined cells corresponds to a source.
+These gadgets correctly realize the model, as shown below: Each sink requires unshaded cells equal to the difference between its current size and 3737, which equals its degree.
 On the other hand, each source supplies a total of two unshaded cells to the adjacent sinks (Figure˜16).
@@
 We construct a reduction via the flow model as illustrated in Figure˜18.
-Each cross-shaped region containing a clue $9$ corresponds to a sink, while each region with a clue $7$ connecting two arms of sinks corresponds to a source.
+Each cross-shaped region containing a clue 99 corresponds to a sink, while each region with a clue 77 connecting two arms of sinks corresponds to a source.
 These sinks and sources function as intended, as we demonstrate below: Each source supplies a total of two unshaded cells to the two connected sinks (Figure˜19).
@@
 Conversely, the interface between two wires of adjacent unit gadgets corresponds to a source.
-These gadgets correctly realize the model, as shown below: The central cell requires a total of $4$ cells from the four wires to form a pentomino, while each wire supplies a specific number of cells to the central cell (Figure˜28).
-Specifically, a wire connected to the board border supplies exactly $1$ cell, effectively adjusting the remaining demand of the central cell to $3$ (when abutting one border) or $2$ (when abutting two borders).
-For the interface between two wires, the sum of their supplies is constrained to $2$; collectively, these two connected wires function as a source.
+These gadgets correctly realize the model, as shown below: The central cell requires a total of 44 cells from the four wires to form a pentomino, while each wire supplies a specific number of cells to the central cell (Figure˜28).
+Specifically, a wire connected to the board border supplies exactly 11 cell, effectively adjusting the remaining demand of the central cell to 33 (when abutting one border) or 22 (when abutting two borders).
+For the interface between two wires, the sum of their supplies is constrained to 22; collectively, these two connected wires function as a source.
 Figure 27: Unit gadget for Five Cells. The undetermined portion is divided into the central cell and four wires. The green cells indicate the connections between the central cell and the wires.
@@
 To simulate the variants of sources, we introduce two specific types of wires (Figure˜29).
-An incoming wire (Figure˜29(a)) can supply either $1$ or $2$ cells, whereas an exclusive wire (Figure˜29(b)) is restricted to supplying either 0 or $2$ cells.
+An incoming wire (Figure˜29(a)) can supply either 11 or 22 cells, whereas an exclusive wire (Figure˜29(b)) is restricted to supplying either 0 or 22 cells.
 A source with exactly one incoming wire corresponds to a biased source oriented toward the incoming side.
-A source with two incoming wires corresponds to a fixed source, as the only pair of flow distributions summing to $2$ is $(1,1)$.
+A source with two incoming wires corresponds to a fixed source, as the only pair of flow distributions summing to 22 is $(1,1)$.
 A source composed of exclusive wires corresponds to an exclusive source.
@@
 6  Conclusion and Future Work
-In this paper, we established the ASP-completeness of RCCP (Theorem˜6) and CCP (Corollary˜7) for planar degree-$3$ bipartite mixed graphs.
-This result is tight in the sense of the maximum degree, as the counting versions of these problems on max-degree-$2$ graphs are obviously polynomial-time solvable.
+In this paper, we established the ASP-completeness of RCCP (Theorem˜6) and CCP (Corollary˜7) for planar degree-33 bipartite mixed graphs.
+This result is tight in the sense of the maximum degree, as the counting versions of these problems on max-degree-22 graphs are obviously polynomial-time solvable.
 However, the complexity for degree-$k$ graphs with $k\geq 4$ remains open.